What are the remaining angles for Triangle of A = 25°, b = 28, a = 18?
sin(25)18=sin(b)28 18sin(b)=28sin(25) sin(b)=28sin(25)18
b=0.71737138

Answer by Alan3354(69443)

What are the remaining angles for Triangle of A = 25°, b = 28, a = 18?
sin(25)18=sin(b)28
-----
sin(25)/18=sin(B)/28
================================


18sin(b)=28sin(25) sin(b)=28sin(25)18
b=0.71737138